1. A pile of 600 sheets of paper has a mass of 2 kg. The pile is 300 mm long, 200 mm wide, 50 mm thick. Calculate:
  2. The thickness of one sheet (2 marks)
  3. The mass of one sheet. (2 marks)
  4. The volume of the pile. (2 marks)
  5. The density of the paper. (2 marks)
1
Expert's answer
2022-03-30T17:43:37-0400

1.

"h_0=\frac cn=83~\mu m,"

2.

"m_0=\frac mn=3.3~g,"

3.

"V=abc=0.003~m^3,"

4.

"\rho=\frac mV=667~\frac{kg}{m^3}."